Convert the following to basic SI units PartA 67 in Express your answer with the appropriate units dValue Units Submit Request Answer Part B

Part B 5.25x106 yr Express your answer with the appropriate units. t= 1 Value Units Submit Request Answer Part C 64 ft/day Express your answer with the appropriate units. Value Units

Part D 4.2x104 mi2 Express your answer with the appropriate units A-Value Units Submit Request Answer Provide Feedback
Answer #1

A.) 1in = 0.0254m

Hence 67in = 1.7018m

B.) 5.25 × 1000000 yrs = 5250000yrs = 5250000×365×24×60×60 seconds = 1.66E14 seconds

C.) 64ft/day = 19.5072/(24×60×60) = 2.258E-4 m/s

D.) 4.2E4 = 1.088 E11 m^2